OpenStudy (anonymous):

A box of 15 parts contains 7 that are defective. A worker picks parts one at a time and attempts to install them. Find the probability of each outcome in​ (a) through​ (d). ​(a) The first two chosen are both good. nothing ​(Round to four decimal places as​ needed.)

OpenStudy (mathmate):

Please check your question, which probably contains typo(s) and incomplete. Cannot have 15 parts of which 77 are defective.

OpenStudy (anonymous):

it a 7

OpenStudy (anonymous):

P(two are good) \[=\frac{C _{2}^{8} }{C_{2}^{15} }\]=?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

ok im so lost 8/15

OpenStudy (anonymous):

and 054

OpenStudy (anonymous):

pl help me out

OpenStudy (anonymous):

im so lost and dum

OpenStudy (mathmate):

(a) Initially, there are 8 good out of 15, so probability of getting the first good one is 8/15. For the next one, there are only 7 good out of 14, so probability is 7/14. For this two step experiment, the probability of getting 2 good out of two is therefore (8/15)*(7/14)=4/15. Alternatively, you could use the formula by @Surjithayer